a helicopter takes vertical take off with an acceleration a=3m/s^2 and u=0m/s. after a certain time t1, a bullet is fired from tthe helicopter . at the point of take off on ground , the sound of the shot is heard at a time t2=30 s after the take off of helicopter . find the velocity of the helicopter  at the moment when the bullet os fired assuming that the velocity of sound is c=320 m/s.

Grade:12

1 Answers

vinamra mahajan
6 years ago
Given, a= 3 m/s^2and u=0.after time t1... v =3t1 ( by v=u +at)Now a bullet is fired V of sound is 320m/s (given)Time taken is 30 seconds (given)Implies that ... S = v*t =9600 m.now... By s=ut +1/2 at^29600= 1/2 * 3 * t^2 6400 = t^2t = 80 sec..Implies that... V= 3 * 80 = 240 m/s